OpenStudy (anonymous):

lim (ln(2+x)-ln2)/(x)= 1/2 x--> 0

OpenStudy (amistre64):

\[\frac{\ln(\frac{2+x}{2})}{x}\]

OpenStudy (amistre64):

\[\ln(1+x)/x = \ln(1+x)^{1/x}\]

OpenStudy (amistre64):

.... typoed it already

OpenStudy (anonymous):

no it is \[ \lim_{x \rightarrow 2} (\ln(2+X)- \ln2)\div x\]

OpenStudy (anonymous):

Just use l'hpital's rule.

OpenStudy (anonymous):

[ln (1 + x/2)] / x multiply and divide by 1/2 \[1/2[\ln (1 + x/2)]/x/2\]

OpenStudy (amistre64):

\[\ln(x+2)-\ln(x) = \ln(\frac{x+2}{x})\]

OpenStudy (anonymous):

now putting x-> 0 it becomes a standard integral of form (ln 1)/0 which is equal to 1, so the ans is 1/2 x 1 = 1/2

OpenStudy (amistre64):

as x-> 2 we get ln(4/2)^(1/2) = ln(sqrt(2))

OpenStudy (anonymous):

amistre i think im right

OpenStudy (anonymous):

wtdu say anwar???

OpenStudy (amistre64):

him; youre right as far as x_.0 perhaps; but the question was amended

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

Why don't use L'hopital's rule? It's going to be easy peasy.

OpenStudy (anonymous):

oh yeah but maybe he wants to show d actual method

OpenStudy (amistre64):

we both get to ln(1+x/2)^(1/x) :)

OpenStudy (anonymous):

no d qstn is ln (2+x) - ln2 whole upon x

OpenStudy (amistre64):

i know; and we both got to ln(1+ x/2)^(1/x)

OpenStudy (anonymous):

does ln 1/0 does not equal 1.

OpenStudy (amistre64):

as x-> 2 we get ln(sqrt(2))

OpenStudy (amistre64):

as x-> 0 we get ln(1)^(.000...0001)

OpenStudy (amistre64):

1^(tiny number) = 1 right?

OpenStudy (amistre64):

1 ^(any#) = 1 lol

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

ln(1) = 0

OpenStudy (anonymous):

This is the definition of the derivative of ln(x) at x = 2. Since the derivative of ln(x) = 1/x, at 2 you get 1/2

OpenStudy (anonymous):

The derivative as x goes to 0 is equal to 1/2

OpenStudy (anonymous):

I think we got the questioner confused :). @lovehap, Did you get what you asked about?

OpenStudy (anonymous):

yes it is \[\lim_{x \rightarrow 2}(\ln(2+x)-\ln(2))/ x\]

OpenStudy (anonymous):

\[f(x)=ln (x),f'(x)=\frac{1}{x}, f'(2)=\lim_{h\to0}\frac{ln(2+h)-ln(2)}{h}=\frac{1}{2}\]

OpenStudy (anonymous):

thank you satellite73!

OpenStudy (amistre64):

Dx(ln(x)) = 1/x f'(2) = 1/2.... yes

OpenStudy (anonymous):

welcome.

OpenStudy (anonymous):

ok thank you!

OpenStudy (kyanthedoodle):

Past-satellite! You'll never believe what the future has in store for you!